OpenStudy (anonymous):

What is the value of x in the equation 8 + x = 3? help plz

OpenStudy (usukidoll):

we need x by itself 8+x =3 so subtract 8 from both sides 8+x = 3 -8 -8 x = 3-8 what's 3-8?